yomomalikebracobama:

2x − 3y − z if x = −2, y = 3, and z = −2

yomomalikebracobama:

Evaluate the following expression using the values given: (5 points) Find 2x − 3y − z if x = −2, y = 3, and z = −2 a −11 b 11 c −15 d 15

yomomalikebracobama:

i dont understand how to solve that all i need is what to do i wanna solve it @tetsxpreme

Ybndrrrip17:

what are you solving for

yomomalikebracobama:

idk

Ybndrrrip17:

Algebra right?

yomomalikebracobama:

yuh

yomomalikebracobama:

how do i solve it -_-

Ybndrrrip17:

how i got my answer i substituted 2x-3x-z with -2 and z with -2 and y with 3 so it equaled up to 2(-2) -3x3-(-2)

Florisalreadytaken:

@yomomalikebracobama wrote:
how do i solve it -_-
\(2x − 3y − z\) if \(x = −2 \wedge y = 3 \wedge z = −2\) just put the values in: \(=2\cdot (-2) − 3\cdot 3 − (-2)\) \(=-4-3\cdot \:3-\left(-2\right)\) \(=-4-9-\left(-2\right)\) \(=-13+2\) \(-11\)

yomomalikebracobama:

so if i never did this befoe and didnt look at the lesson explain it like i am just learnig it

Ybndrrrip17:

you can do either way they both lead to the right answer

yomomalikebracobama:

not that ididnt watch the lesson....

ILOVESPAGHETTI:

Well in this math problem, you just replace the variables with the actual numbers then solve like these lovely people did!

Ybndrrrip17:

basically make it seem like you know what your doing but also like your new to the lesson

ShadowOfDeath:

I believe what you need to do to solve this problem is to plug each of the numbers into its designated places. Ex: Find 3x-4y-z, if x=2, y=3, and z=4. First, you need to find out what the first term would equal. Plug the 2 into where the x is and then multiply 2 by 3. 3(2)=3 x 2=6 Next we need to find out what the y is. Do the same as you did for x. Plug 3 into where the y is then multiply. 4(3)=4 x 3=12 Lastly, we come to the final term. Once again, we need to plug in what number z is equal to. In this case, it is 2. So the final answer to this would be 6-12-2, which, in solving this, would be -8 because 6-12=-6. -6-2 is -8 because of the fact that a negative plus a negative would still be negative.

ShadowOfDeath:

Do this with this problem and you'll have your answer!

yomomalikebracobama:

ok